Find the maclaurin series expansion of f(x)= 1/1+x hence or otherwise find the expansion of 1/1+x^2 by integrating both sides of your result, find the series expansion of arctanx, by also putting x=1 find the expression Ο€/4 hence show that Ο€=3.142


1
Expert's answer
2021-11-11T12:16:12-0500

Maclaurin series expansion


"\dfrac{1}{1-t}=\displaystyle\sum_{k=0}^\infin t^k, -1<t<1"

Then


"\dfrac{1}{1+t^2}=\dfrac{1}{1-(-t^2)}=\displaystyle\sum_{k=0}^\infin (-t^2)^k"

"=\displaystyle\sum_{k=0}^\infin (-1)^kt^{2k}, -1<t<1"

Integrate both sides


"\displaystyle\int_{0}^x\dfrac{1}{1+t^2}dt=\displaystyle\int_{0}^x\displaystyle\sum_{k=0}^\infin (-1)^kt^{2k}dx"

"\arctan x=\displaystyle\sum_{k=0}^\infin \dfrac{(-1)^kx^{2k+1}}{2k+1}, -1\leq x\leq1"

"\arctan 1=\displaystyle\sum_{k=0}^\infin \dfrac{(-1)^k(1)^{2k+1}}{2k+1}""\dfrac{\pi}{4}=\arctan 1=\displaystyle\sum_{k=0}^\infin \dfrac{(-1)^k}{2k+1}"

"=1-\dfrac{1}{3}+\dfrac{1}{5}-\dfrac{1}{7}+\dfrac{1}{9}-\dfrac{1}{11}+..."

"\pi=4(1-\dfrac{1}{3}+\dfrac{1}{5}-\dfrac{1}{7}+\dfrac{1}{9}-\dfrac{1}{11}+...)"

"\pi\approx3.1416"